OpenStudy (anonymous):

Bruce has a bottle that contains 60% of lemon juice and the rest water. The bottle has 1 liter of water. Part A: Write an equation in one variable that can be used to find the total number of liters of lemon juice and water in the bottle. Define the variable used in the equation. Part B: How many liters of lemon juice are present in the bottle? Show your work.

OpenStudy (anonymous):

\(\Huge{Please~Help!!!!}\)

OpenStudy (anonymous):

\(Do~you~know~how~to~solve~it?\)

OpenStudy (ghostgate):

I know how to solve Part B, I'm not entirely sure I can solve Part A.

OpenStudy (anonymous):

me too

OpenStudy (ghostgate):

So the problem is you need help with Part A?

OpenStudy (anonymous):

Yes. I don't know how to write it in a one-variable equation

OpenStudy (ghostgate):

Hold on, I'll see what I can come up with.

OpenStudy (anonymous):

Thank you.

OpenStudy (ghostgate):

I suppose you may want to write the equation like this: 1(0.4) + x(0.6) = 1 Due to how if you multiply the liters times the decimal percentage you would equal the decimal equivalent of 1.

OpenStudy (anonymous):

But what would x represent?

OpenStudy (ghostgate):

x would represent the amount of liters that the lemon juice would have. Let me see if I try to solve it: 0.4 + 0.6x = 1 0.6x = 0.4 Nevermind... hold on.

OpenStudy (anonymous):

okay

OpenStudy (ghostgate):

Wait... 1/0.4 - 0.6(2.5) = x? With x representing the lemon juice amount.

OpenStudy (ghostgate):

As you have probably noticed, I don't entirely know what I am doing. I'll keep trying though.

OpenStudy (anonymous):

thank you so much

OpenStudy (ghostgate):

I think I know it, you would probably write it like this: x/0.6 = 2.5 x represents the liters of lemon juice.

OpenStudy (ghostgate):

Because 2.5 x 0.6 = 1.5 which is the amount of liters used.

OpenStudy (anonymous):

Thank you ghost for your time and effort you put towards helping me. I am eternally grateful.

OpenStudy (ghostgate):

You're welcome! I'm sorry if the answer isn't correct.